# Conan: A Chunkwise Online Network for Zero-Shot Adaptive Voice Conversion

### v2

@@ III-C Adaptive Style Encoder
To comprehensively capture and transfer styles from reference speech, we introduce the Adaptive Style Encoder. As shown in Figure 2, reference mel-spectrogram $m_{rf}$ is initially refined through convolutional blocks. Subsequently, a downsampling layer pools the output into chunk-level representations. In our experiments, we set the chunk size to 80 ms to effectively incorporate emotion and prosody. Then, convolutional stacks capture correlations at the chunk level.
Next, we employ a linear projection to map the outputs into a low-dimensional latent space for efficient code index lookup [37]. The Clustering Vector Quantization (CVQ) layer [38] extracts style representations from these latent inputs, establishing an information bottleneck that effectively filters out non-style information [39]. Through dimensionality reduction via linear projection and the bottleneck mechanism of CVQ, we decouple speech styles from timbre and content.
Compared to traditional VQ [40], CVQ employs a dynamic initialization strategy during training. This ensures that less frequently or unused code vectors are updated more substantially than frequently used ones, effectively addressing the codebook collapse issue [38]. We train the CVQ layer using a combination of the VQ loss and contrastive loss, defined as:
$\mathcal{L}_{CVQ}=$ $\|sg[z_{e}(x)]-e\|^{2}_{2}+\beta\|z_{e}(x)-sg[e]\|^{2}_{2}+\mathcal{L}_{Contras},$ (7)
where $z_{e}(x)$ is the encoder output for input $x$, $e$ is the selected code from the codebook, $\text{sg}(\cdot)$ is the stop-gradient operator, and $\beta$ is commitment loss. The contrastive loss is:
$\mathcal{L}_{Contras}=-\log\frac{e^{\text{sim}(e_{k},\hat{z}_{i}^{+})/\tau}}{\sum*{i=1}^{N}e^{\text{sim}(e_{k},\hat{z}_{i}^{-})/\tau}}.$ (8)
For each code $e_{k}$, we directly select the closest $\hat{z}_{i}^{+}$ as the positive pair and sample other farther $\hat{z}_{i}^{-}$ as negative pairs.
Upon generating the style embedding, it is crucial to adaptively align this embedding with the sum $z_{ct-i}$ of $z_{c-i}$ and $z_{t}$ [41]. For this purpose, we introduce the Align Attention module, incorporating the Scaled Dot-Product Attention mechanism [42]. Before inputting the detailed style embedding into the attention module, we include positional encoding embeddings. Within this module, $z_{ct-i}$ serves as the query, while the style embedding functions as both key and value. Ultimately, this process yields aligned style $z_{s-i}$, effectively facilitating the modeling and transfer of reference styles.
@@ III-D Causal Shuffle Vocoder
The chunkwise processing in online VC makes it incompatible with offline VC models and introduces boundary artifacts. Enforcing causality through autoregressive methods increases latency, while using zero-padding in non-causal vocoders leads to spectral artifacts. To address this, after generating the mel-spectrogram $m_{gn-i}$ using a causal mel decoder composed of causal convolutions, we build a fully causal streaming vocoder. Based on high-quality HiFi-GAN [43], we replace standard convolutions with causal convolutions and implement upsampling via pixel shuffle, resulting in a strictly causal vocoder with high audio quality, named Causal Shuffle Vocoder.
As illustrated in Figure 1 (c), $m_{gn-i}$ first passes through a single causal convolutional layer, producing an initial feature map of shape $(C,T_{0})$. The network then comprises $N$ upsampling stages. At the $n$-th stage, a causal convolution projects the channel dimension from $C$ up to $r_{n}C$, yielding an intermediate feature map of shape $(r_{n}C,T_{n-1})$. Next, pixel shuffling rearranges these $r_{n}C$ channels into a new tensor of shape $(C,r_{n}T_{n-1})$, $r_{n}T_{n-1}=T_{n}$, thereby performing temporal upsampling. Because pixel shuffling only reorders channels without introducing information from future frames, the network’s output at time step $t$ depends solely on input frames from time $1$ through $t$. By choosing the upsampling factors, the final temporal resolution becomes $T_{0}\prod_{n=1}^{N}r_{n}$, recovering the original audio sampling rate.
@@
By combining causal convolution with pixel shuffling, the waveform produced at each frame depends exclusively on past mel-spectrogram samples. Causal convolutions pad solely on the left, avoiding spectral distortions from symmetric or right‐side padding. Moreover, pixel shuffling upscales by rearranging channels into a finer temporal grid without waiting for future frames, thus eliminating checkerboard artifacts. This approach matches HiFi‐GAN quality in an online, low‐latency setting because each step is a single forward pass, avoiding cumulative delay. Moreover, since pixel shuffling uniformly distributes low-resolution features along the time axis, the model can more effectively reconstruct audio details, yielding clear, natural speech even under strict causal constraints.
@@ III-E Training and Inference Procedures
Training Procedure. First, for Stream Content Extractor, we use the cross-entropy between the predicted content label and the label extracted by HuBERT. During training of the main model, we adopt the multi-length discriminator [44] to maintain perceptual quality. The total loss comprises $\mathcal{L}_{\mathrm{CVQ}}$ for the adaptive style encoder; $\mathcal{L}_{\mathrm{pitch}}$, the MSE between the ground truth and the predicted $F_{0}$; $\mathcal{L}_{\mathrm{mae}}$ and $\mathcal{L}_{\mathrm{ssim}}$, the MAE and SSIM losses between the predicted and ground truth mel‐spectrograms; and $\mathcal{L}_{\mathrm{GAN}}$, the LSGAN‐style adversarial loss [45], whose objective is to minimize the distribution distance between predicted and ground truth mel‐spectrograms. Finally, the loss for Causal Shuffle Vocoder follows HiFiGAN [43], including the GAN loss, mel loss, and feature‐map loss.
Inference Procedure. As illustrated in Figure 3, we first feed the entire reference speech into the model to provide timbre and stylistic information. During chunk‐wise online inference, we wait until the input reaches a predefined chunk size before passing it to the model, because our generation speed for each chunk is faster than the chunk’s duration, online generation becomes possible. To ensure temporal continuity, we employ a sliding context‐window strategy. Concretely, at each generation step, we not only input the source speech corresponding to the current chunk but also include the preceding context within the model’s receptive field. From the model’s output, we then extract only the segment required for this chunk, thus preserving causality. As a result, overlapping portions generated across successive chunks remain identical, yielding perfectly smooth boundaries between concatenated segments.
@@ IV-A Experimental Setup
Dataset. We use the LibriTTS dataset [46] for our experiments, which contains 586 hours of English speech from 2,456 speakers. All speech is downsampled to 16 kHz. To evaluate Conan’s performance in zero‐shot scenarios, we generalized the source model to the VCTK dataset, which comprises 44 hours of clean speech from 110 English speakers. For a fair comparison, we reference the StreamVC [3] data‐partitioning strategy. We use the test‐clean subset of LibriTTS as the source speech and then randomly select 6 unseen speakers from the VCTK dataset to form the out‐of‐domain test set.
Implementation Details. We set the sample rate to 16 kHz, the window size to 1024, the hop size to 320, and the number of mel bins to 80 to derive mel-spectrograms from raw waveforms. We set the chunk size to 80 ms and employ a 6-layer Emformer with 2 right context chunks as the full setting. For the fastest setting, the chunk size is 20 ms, and a 3-layer Emformer without right context chunks is employed. The default size of the codebook for CVQ is 128.
@@
Evaluation Details. We objectively and subjectively evaluate our model based on three criteria: real‐time content accuracy, conversion quality, and speaker similarity. Content accuracy is measured by Word Error Rate (WER) and Character Error Rate (CER) using a HuBERT‐Large ASR model  [^1] https://github.com/resemble-ai/Resemblyzer. Speaker similarity (SIM) is assessed with Resemblyzer  [^2] https://huggingface.co/facebook/hubert-large-ls960-ft. For subjective metrics, we conduct mean opinion score (MOS) and comparative mean opinion score (CMOS) evaluations. Both metrics are rated on a 1–5 scale and reported with 95% confidence intervals. Specifically, we use MOS‐Q to judge conversion quality (for clarity and naturalness) and MOS‐S to assess speaker similarity (for timbre and styles) [47]. For CMOS-Q and CMOS-S, listeners are tasked with comparing pairs of speech produced by different systems and expressing their preferences. All participants are fairly compensated and agree that the results will be used for scientific research.
@@ IV-B Main Results
TABLE I: Evaluation results of Conan in comparison with baseline approaches. We conduct both subjective and objective evaluations.
Method Method Content Accuracy Speaker Similarity Conversion Quality
@@
Conan (Fastest) 10.72% 4.74% 78.26% 3.85 $\pm$ 0.13 3.99 $\pm$ 0.11
We use several popular voice‐conversion models, including VQMIVC [4], BNE‐PPG‐VC [48], Diff‐VCTK [49], QuickVC [50], and StreamVC [3]. Since StreamVC is not open source, to ensure a fair comparison, we adopted the same dataset split and objective evaluation protocols they used. We rely on the objective metrics reported in their paper and perform subjective assessments based on their demo pages. Specifically, we collect utterances from speakers in the test-clean subset of LibriTTS as the source speech, and randomly select 6 speakers from the VCTK dataset as the reference speech sample.
As shown in Table I, Conan in the full setting delivers the highest speaker similarity (SIM, MOS-S), indicating that the Adaptive Style Encoder effectively models the speaker’s style and therefore achieves superior zero-shot similarity. Conan also attains the best conversion quality (MOS-Q), demonstrating that the streaming framework enhances both synthesis quality and naturalness. For content accuracy, Conan records the lowest WER, while its CER is slightly higher than that of StreamVC. This is mainly because StreamVC does not attempt style transfer and instead reuses the source pitch and energy distributions, giving it a CER advantage. Finally, the metrics for Source (Vocoder) and Reference (Vocoder) are both excellent, further highlighting the outstanding online reconstruction capability of the Causal Shuffle Vocoder.
TABLE II: Latency and RTF for Conan’s full and fastest settings tested on a single A100 80 GB GPU. Latency includes the sum of each module’s delay, plus the chunk size, and the right context.
Method Full Setting Fastest Setting
@@
As shown in Table II, in the faster setting, Conan uses a 20 ms chunk size to perform strictly causal, chunk‐wise online voice conversion with only 37 ms of latency, yet still achieves high performance metrics. Under Conan’s full setting with 80 ms chunk size and 2 right context chunks, we achieve substantially higher quality, nearly matching offline VC performance, while still meeting real‐time latency constraints.
@@ IV-C Ablation Study
As shown in Table III, we conduct ablation studies to demonstrate the effectiveness of various design choices within Conan. When the original HuBERT is used instead of the Stream Content Encoder, we observe a substantial deterioration in both synthesis quality and WER, underscoring the incompatibility of HuBERT with streaming scenarios. Furthermore, when right‐context chunks are omitted, both MOS‐Q and WER worsen, indicating that the size of the lookahead can be adjusted as a trade‐off between efficiency and quality. Furthermore, removing the Adaptive Style Encoder results in a marked decrease in MOS-S, confirming that our approach effectively captures speaker styles. Finally, substituting the Causal Shuffle Vocoder with a zero-padded causal HiFi-GAN leads to a decline in MOS-Q, as the zero padding introduces checkerboard artifacts, causing a decline in synthesis quality.
@@ V Conclusion
In this paper, we introduce Conan, a method that converts speech by preserving the original content while adopting the timbre and styles of a reference utterance. At its core, Conan employs the Streaming Content Extractor based on Emformer to encode source audio with minimal latency, ensuring semantic information is captured in real time. To adapt to styles of reference speech, Adaptive Style Encoder analyzes the reference speech and derives detailed stylistic representations that guide the conversion process. Finally, Causal Shuffle Vocoder, built upon a fully causal HiFiGAN architecture enhanced with pixel-shuffle operations, reconstructs high-fidelity audio with low generation delay. Experimental results indicate that Conan surpasses existing baseline approaches, achieving higher scores in both subjective and objective metrics.

### v3

@@ III-C Adaptive Style Encoder
To comprehensively capture and transfer speaker identity from reference speech, we introduce the Adaptive Style Encoder. As shown in Figure 2, reference mel-spectrogram $m_{rf}$ is initially refined through convolutional blocks. Subsequently, a downsampling layer pools the output into chunk-level representations. In our experiments, we set the chunk size to 80 ms to incorporate styles, which may include emotion and prosody.
Next, we employ a linear projection to map the outputs $z$ into a low-dimensional latent space for efficient code index lookup [37]. The Clustering Vector Quantization (CVQ) layer [38] extracts style representations from these latent inputs, establishing an information bottleneck that effectively filters out non-style information [39]. By applying linear projection and the CVQ mechanism, we capture fine-grained speaker identity features except global timbre and content [31].
Compared to traditional VQ [40], CVQ employs a dynamic initialization strategy during training. This ensures that less frequently or unused code vectors are also updated as frequently used ones do, effectively addressing the codebook collapse issue [38]. We train the CVQ layer using a combination of the VQ loss and contrastive loss, defined as:
$\mathcal{L}_{CVQ}=$ $\|sg[z]-e\|^{2}_{2}+\beta\|z-sg[e]\|^{2}_{2}+\mathcal{L}_{Contras},$ (7)
[…]
$\mathcal{L}_{Contras}=-\log\frac{e^{\text{sim}(e,z^{+})}}{e^{\text{sim}(e,z^{+})}+\sum_{i=1}^{N}e^{\text{sim}(e,z_{i}^{-})}}.$ (8)
After generating the style embedding, we adaptively align it with the content, taking into account the influence of global timbre on style selection, by matching it to $z_{ct-i}$, the concatenation of $z_{c-i}$ and $z_{t}$ [41]. For this purpose, we introduce the Align Attention module, incorporating the Scaled Dot-Product Attention mechanism [42]. Before inputting the detailed style embedding into the attention module, we also add positional encoding embeddings. Within this module, $z_{ct-i}$ serves as the query, while the style embedding functions as both key and value. Ultimately, we get the detailed speaker identity $z_{s-i}$.
@@ III-D Causal Shuffle Vocoder
Enforcing online VC causality through frame-by-frame autoregressive methods degrades the quality of early generated segments and incurs extra computational overhead, while using zero-padding in vocoders leads to spectral artifacts. To address this, after generating the mel-spectrogram $m_{gn-i}$ using a causal mel decoder composed of causal convolutions, we build Causal Shuffle Vocoder. Based on high-quality HiFi-GAN [43], we replace standard convolutions with causal convolutions and implement upsampling via pixel shuffle, resulting in a strictly causal vocoder with high audio quality.
As illustrated in Figure 1 (c), $m_{gn-i}$ first passes through a single causal convolutional layer, producing an initial feature map of shape $(C,T_{0})$, where $C$ and $T_{0}$ represent the channel and time dimensionality, respectively. The network then comprises $N$ upsampling stages. At the $n$-th stage, a causal convolution projects the channel dimension from $C$ up to $r_{n}C$, yielding an intermediate feature map of shape $(r_{n}C,T_{n-1})$. Next, pixel shuffling rearranges these $r_{n}C$ channels into a new tensor of shape $(C,r_{n}T_{n-1})$, where $r_{n}T_{n-1}=T_{n}$, to achieve temporal upsampling. Because pixel shuffling only reorders channels without introducing information from future frames, the network’s output is frame-level causal. By choosing the upsampling factors, the final temporal resolution becomes $T_{0}\prod_{n=1}^{N}r_{n}$, recovering the original audio sampling rate.
@@
By combining causal convolution with pixel shuffling, the waveform produced at each frame depends exclusively on past mel-spectrogram samples. Causal convolutions pad solely on the left, avoiding spectral distortions from symmetric or right‐side padding. Pixel shuffling achieves temporal upsampling by simply reorganizing feature‑map channels into a finer time grid, rather than by learning interpolation kernels that span future frames. Because it never blends or spreads information unevenly across time, it inherently avoids the checkerboard artifacts typical of transposed convolutions [44]. In our strictly causal setup, this means we can match the audio fidelity of HiFi‑GAN while operating online with minimal latency.
@@ III-E Training and Inference Procedures
Training Procedure. First, for Stream Content Extractor, we use the cross-entropy between the predicted content label and the label extracted by HuBERT. Then, the total loss for training the main model comprises $\mathcal{L}_{\mathrm{CVQ}}$ for the adaptive style encoder; $\mathcal{L}_{\mathrm{pitch}}$, the MSE between the ground truth and the predicted $F_{0}$; $\mathcal{L}_{\mathrm{mae}}$ and $\mathcal{L}_{\mathrm{ssim}}$, the MAE and SSIM losses between the predicted and ground truth mel‐spectrograms; and $\mathcal{L}_{\mathrm{GAN}}$, the LSGAN‐style adversarial loss [45], whose objective is to minimize the distribution distance between predicted and ground truth mel‐spectrograms. Finally, the loss for Causal Shuffle Vocoder follows HiFiGAN [43], including the GAN loss, mel loss, and feature‐map loss.
Inference Procedure. As shown in Figure 3, we first feed the entire reference speech into the model to provide timbre and stylistic information. During chunk‑wise online inference, we wait until the input reaches a predefined chunk size before passing it to the model. Because our generation speed for each chunk is faster than the chunk’s duration, online generation becomes possible. To ensure temporal continuity, we employ a sliding context window strategy. At each generation step, we not only input the source speech of the current chunk but also include the preceding context. From the model’s output, we extract only the segment for this chunk. As the context covers the receptive field, consistent overlapping segments can be generated, ensuring smooth transitions at chunk boundaries.
@@ IV-A Experimental Setup
Dataset. We randomly use the source and reference speakers in the train subset of the LibriTTS dataset [46], which contains 586 hours of English speech from 2,456 speakers. All speech is downsampled to 16 kHz. To evaluate Conan’s performance in zero‐shot scenarios, we include the VCTK dataset, which comprises 44 hours of clean speech from 110 English speakers. For a fair comparison, we reference the StreamVC [3] data‐partitioning strategy. We use the test‐clean subset of LibriTTS as the source speech and then randomly select 6 unseen speakers from the VCTK dataset to form the test set.
Implementation Details. We set the sample rate to 16 kHz, the window size to 1024, the hop size to 320, and the number of mel bins to 80 to derive mel-spectrograms from raw waveforms. We set the chunk size to 80 ms and employ a 6-layer Emformer with 2 right context chunks as the full setting. For the fast setting, the chunk size is 20 ms, and a 3-layer Emformer without right context chunks is employed. The default size of the codebook for CVQ is 128.
@@
Evaluation Details. We objectively and subjectively evaluate our model based on three criteria: real‐time content accuracy, quality, and speaker similarity. Content accuracy is measured by Word Error Rate (WER) and Character Error Rate (CER) using a HuBERT‐Large ASR model  [^1] https://github.com/resemble-ai/Resemblyzer. Speaker similarity (SIM) is assessed with Resemblyzer  [^2] https://huggingface.co/facebook/hubert-large-ls960-ft. For subjective metrics, we conduct mean opinion score (MOS) and comparative mean opinion score (CMOS) evaluations. Both metrics are rated on a 1–5 scale and reported with 95% confidence intervals. Specifically, we use MOS‑Q to judge quality (clarity and naturalness) and MOS‑S to assess speaker similarity (timbre and styles) [47]. For CMOS‑Q and CMOS‑S, listeners compare pairs of speech samples produced by different systems. We randomly selected at least 20 pairs with more than 6 unseen speakers for each subjective evaluation, with at least 15 listeners per pair. All participants were fairly compensated and provided informed consent for the use of their responses in scientific research.
@@ IV-B Main Results
TABLE I: Evaluation results of Conan and baseline models, including both objective (WER, CER, SIM) and subjective (MOS-S, MOS-Q) evaluations. All baselines, except for StreamVC, are evaluated in an offline setting.
Method Method Content Accuracy Speaker Similarity Quality
@@
Conan (Fast) 10.72% 4.74% 78.26% 3.85 $\pm$ 0.13 3.99 $\pm$ 0.11
We use several popular voice‐conversion models, including VQMIVC [4], BNE‐PPG‐VC [48], Diff‐VCTK [49], QuickVC [50], and StreamVC [3] as baselines. Since StreamVC is not open-sourced, to ensure a fair comparison, we adopt the same dataset split and objective evaluation protocols that they used. We rely on the objective metrics reported in their paper and perform subjective assessments based on their demo pages.
As shown in Table I, Conan in the full setting delivers the highest speaker similarity (SIM, MOS-S), indicating that the Adaptive Style Encoder and timbre encoder effectively model the speaker identity and therefore achieve superior zero-shot similarity. Conan also attains the best quality (MOS-Q), demonstrating that the streaming framework enhances synthesis quality. For content accuracy, Conan records the lowest WER, while its CER is slightly higher than that of StreamVC. This is mainly because StreamVC reuses the source pitch and energy values, giving it a CER advantage. Finally, the metrics for Source (Vocoder) and Reference (Vocoder) are both excellent, further highlighting the outstanding online reconstruction capability of the Causal Shuffle Vocoder.
TABLE II: Latency and RTF for Conan’s full and fast settings tested on a single A100 80 GB GPU. Latency includes the sum of each module’s delay, plus the chunk size, and the right context.
Method Full Setting Fast Setting
@@
As shown in Table II, in the faster setting, Conan uses a 20 ms chunk size to perform strictly causal, chunk‐wise online voice conversion with only 37 ms of latency, yet still achieves high performance metrics. Under Conan’s full setting with 80 ms chunk size and 2 right context chunks, we achieve high quality while still meeting real‐time latency constraints.
@@ IV-C Ablation Study
As shown in Table III, we conduct ablation studies to demonstrate the effectiveness of various design choices within the full setting. When the original HuBERT is used instead of the Stream Content Encoder, we observe a substantial deterioration in both synthesis quality and WER, underscoring the incompatibility of HuBERT with streaming scenarios. Furthermore, when right‐context chunks are omitted, all metrics worsen, indicating that the size of the lookahead can be adjusted as a trade‐off between efficiency and quality. Furthermore, removing the Adaptive Style Encoder results in a marked decrease in MOS-S, confirming that our approach effectively captures speaker identity. Finally, substituting the Causal Shuffle Vocoder with a zero-padding causal HiFi-GAN leads to a decline in MOS-Q, as the zero padding causal HiFi-GAN causes a decline in synthesis quality.
@@ V Conclusion
In this paper, we introduce Conan, a streaming voice conversion method that preserves the linguistic content of a source utterance while adopting the speaker information of a reference utterance. At its core, Conan employs the Streaming Content Extractor based on Emformer to encode source audio with minimal latency, ensuring semantic information is captured in real time. To transfer the reference speaker identity, Adaptive Style Encoder analyzes the reference speech and derives detailed speaker representations that guide the conversion process. Finally, Causal Shuffle Vocoder, built upon a fully causal HiFiGAN architecture enhanced with pixel-shuffle operations, reconstructs high-fidelity audio with low generation delay. Experimental results show that Conan surpasses existing baseline approaches, achieving better performance according to both subjective and objective metrics.